The time taken to pluck the tea leaves in a certain plantation has a normal distribution with mean of 25 hours per week and a standard deviation of 4 hours. Calculate the probability that the tea leaves can be plucked at this plantation in the following period of time:


a) More than 30 hours

[3 marks]


b) Between 18 and 34 hours

[4 marks]


c) Between 25 and 34 hours

[4 marks]




1
Expert's answer
2021-02-02T04:37:04-0500

"\mu= 25,~\sigma=4,"

"Z=\frac{t-\mu}{\sigma},"

1)

"P(t>30)=1-P(Z\leq\frac{30-25}{4})=\\ 1-P(Z\leq1.25)=1-0.8944=0.1056,"

2)

"P(18<t<34)=\nP(Z\u2264 \n\\frac{\n34\u221225}{4}\n\u200b\t\n )\u2212P(Z\u2264 \n\\frac{\n18\u221225}{4}\n\u200b\t\n )=P(Z\u2264 \n2.25)\u2212P(Z\u2264 \n-1.75)=\n0.9878\u22120.0401=0.9477,"

3)

"P(25<t<34)=\nP(Z\u2264 \n\\frac{\n34\u221225}{4}\n\u200b\t\n )\u2212P(Z\u2264 \n\\frac{\n25\u221225}{4}\n\u200b\t\n )=P(Z\u2264 \n2.25 )\u2212P(Z\u2264 \n0)=\n0.9878\u22120.5=0.4878."